# HC-Mamba: Vision MAMBA with Hybrid Convolutional Techniques for Medical Image Segmentation

### v1

@@ Abstract
Automatic medical image segmentation technology has the potential to expedite pathological diagnoses, thereby enhancing the efficiency of patient care. However, medical images often have complex textures and structures, and the models often face the problem of reduced image resolution and information loss due to downsampling. To address this issue, we propose HC-Mamba, a new medical image segmentation model based on the modern state space model Mamba. Specifically, we introduce the technique of dilated convolution in the HC-Mamba model to capture a more extensive range of contextual information without increasing the computational cost by extending the perceptual field of the convolution kernel. In addition, the HC-Mamba model employs depthwise separable convolutions, significantly reducing the number of parameters and the computational power of the model. By combining dilated convolution and depthwise separable convolutions, HC-Mamba is able to process large-scale medical image data at a much lower computational cost while maintaining a high level of performance. We conduct comprehensive experiments on segmentation tasks including skin lesion, and conduct extensive experiments on ISIC17 and ISIC18 to demonstrate the potential of the HC-Mamba model in medical image segmentation. The experimental results show that HC-Mamba exhibits competitive performance on all these datasets, thereby proving its effectiveness and usefulness in medical image segmentation.
@@ 1 Introduction
We conducted comprehensive experiments on segmentation tasks including organ segmentation, skin lesion and brain tumor images, and conduct extensive experiments on ISIC17 and ISIC18[18] to demonstrate the potential of the HC-Mamba model in medical image segmentation. The experimental results show that HC-Mamba exhibits competitive performance on all these datasets, thereby proving its effectiveness and usefulness in medical image segmentation.
@@ 3.1 Datasets
We conduct comprehensive experiments on HC-Mamba for medical image segmentation tasks. Specifically, we evaluate the performance of HC-Mamba on medical image segmentation tasks on the ISIC17, ISIC18 datasets.
@@ 3.2 Results
We compare HC-Mamba with some state-of-the-art models and some recent mamba-based model, presenting the experimental results in Table 1. In order to fully demonstrate that HC-Mamba’s potential in medical image segmentation tasks directly benefits from SSM, we did not use any pre-training strategies.
For the ISIC2017 and ISIC2018 datasets, HC-Mamba performs well on mIoU and Dice compared to other models. Specifically, HC-Mamba has a 1.46% and 1% advantage over MedMamba on mIoU and Dice, respectively, while it has a 2.74% and 1.7% advantage over Unet on mIoU and Dice, respectively.

### v2

@@ Abstract
Automatic medical image segmentation technology has the potential to expedite pathological diagnoses, thereby enhancing the efficiency of patient care. However, medical images often have complex textures and structures, and the models often face the problem of reduced image resolution and information loss due to downsampling. To address this issue, we propose HC-Mamba, a new medical image segmentation model based on the modern state space model Mamba. Specifically, we introduce the technique of dilated convolution in the HC-Mamba model to capture a more extensive range of contextual information without increasing the computational cost by extending the perceptual field of the convolution kernel. In addition, the HC-Mamba model employs depthwise separable convolutions, significantly reducing the number of parameters and the computational power of the model. By combining dilated convolution and depthwise separable convolutions, HC-Mamba is able to process large-scale medical image data at a much lower computational cost while maintaining a high level of performance. We conduct comprehensive experiments on segmentation tasks including organ segmentation and skin lesion, and conduct extensive experiments on Synapse , ISIC17 and ISIC18 to demonstrate the potential of the HC-Mamba model in medical image segmentation. The experimental results show that HC-Mamba exhibits competitive performance on all these datasets, thereby proving its effectiveness and usefulness in medical image segmentation.
@@ 1 Introduction
We conducted comprehensive experiments on segmentation tasks including organ segmentation and skin lesion , and conduct extensive experiments on Synapse, ISIC17 and ISIC18[18] to demonstrate the potential of the HC-Mamba model in medical image segmentation. The experimental results show that HC-Mamba exhibits competitive performance on all these datasets, thereby proving its effectiveness and usefulness in medical image segmentation.
@@ 3.1 Datasets
We conduct comprehensive experiments on HC-Mamba for medical image segmentation tasks. Specifically, we evaluate the performance of HC-Mamba on medical image segmentation tasks on the Synapse, ISIC17 and ISIC18 datasets.
@@
  • Synapse:The Synapse Multi-Organ Segmentation dataset contains CT images from over 30 patients for abdominal multi-organ segmentation. This dataset includes segmentation labels for organs such as the liver, spleen, and pancreas, totaling over 4200 images. The images are of uniform resolution and provide accurate labeling of organ contours to support the needs of medical image processing and deep learning research. The dataset is widely used for training and evaluation of medical image segmentation algorithms, especially for improving the accuracy and efficiency of automatic segmentation.
@@ 3.2 Results
We compare HC-Mamba with some state-of-the-art models and some recent mamba-based model, presenting the experimental results in Table 1 and 2. In order to fully demonstrate that HC-Mamba’s potential in medical image segmentation tasks directly benefits from SSM, we did not use any pre-training strategies.
For the ISIC2017 and ISIC2018 datasets, HC-Mamba performs well on mIoU and Dice compared to other models. Specifically, HC-Mamba has a 1.46% and 1% advantage over MedMamba on mIoU and Dice, respectively, while it has a 2.74% and 1.7% advantage over Unet on mIoU and Dice.
@@
For the Synapse datasets, HC-Mamba performs well on DSC and HD95 compared to other models. Specifically, HC-Mamba has a 0.89% and 2.48% advantage over MedMamba and VM-Unet on Dice.
Table 2: Comparative experimental results on the Synapse dataset. DSC of every single class is also reported. (Bold indicates the best.)
[…]
V-Net [26] 68.81 - 75.34 51.87 77.10 80.75 87.84 40.05 80.56 56.98
DARR [27] 69.77 - 74.74 53.77 72.31 73.24 94.08 54.18 89.90 45.96
UNet [6] 76.85 42.70 89.07 69.72 77.77 68.60 93.43 53.98 86.67 75.58
Att-UNet [24] 77.77 39.02 89.55 68.88 77.98 71.11 93.57 58.04 87.30 75.75
TransUnet [7] 76.38 34.69 86.13 62.03 80.77 75.92 92.98 54.76 83.98 74.52
Swin U-Net [10] 77.15 29.55 83.49 64.55 81.30 77.63 92.31 54.60 88.68 74.62
VM-UNet[21] 79.08 32.21 84.40 67.41 84.16 80.74 92.07 56.90 87.51 79.42
MedMamba[22] 80.67 28.15 88.23 69.53 82.22 79.49 96.89 58.57 92.29 79.12
Hc-Mamba 81.56 26.32 90.92 69.65 85.57 79.27 97.38 54.08 93.49 80.14